OpenStudy (anonymous):

A small toy company makes two types of toys: dolls and toy cars. The expense to make a toy car is $5. The expense to make a doll is $7. The company spends $780 a day making toys. They also make 20 toy cars fewer than twice the number of dolls they make. Construct a system of equations that describes this situation.

OpenStudy (anonymous):

@jhonyy9 Wanna help me?

jhonyy9 (jhonyy9):

like idea note toy car with x and doll car with y than rewrite your exercise ok ?

OpenStudy (anonymous):

i dont understand...

jhonyy9 (jhonyy9):

a toy car =x a doll car =y so than yor exersice will be for an x need expend 5 dollars for an y need expend 7 dollars total spends 780 dollars a day to making toys

jhonyy9 (jhonyy9):

so the next frase i dont understand exactly

jhonyy9 (jhonyy9):

20 x fewer than twice of y

jhonyy9 (jhonyy9):

how is this 20x fewr than twice of y ?

jhonyy9 (jhonyy9):

fewer - sorry

jhonyy9 (jhonyy9):

sorry but i think that not can understanding it exactly correct

OpenStudy (anonymous):

It's fine.

jhonyy9 (jhonyy9):

sorry this mean that the company making with 20 toy cars in minus than twice of dolls is it correct ?

jhonyy9 (jhonyy9):

780/5 =156 so this mean 156 toys per day but make 20 toy cars in minus than twice of dolls so this mean 156-20=136 so thia 136 is twice of dolls so than 136/2 =68 dolls per day